Find the odds in favor of the indicated event. Spinning an A on the spinner pictured below. (The sectors are of equal size.) AIA O 2 to 1 O 3to5 O 1 to 3 O 3 to 1
Answer #1

Total sectors of circle = 8

total numbers of A = 3

therefore number of sectors not having A = 8 - 3 = 5

Odds in favour = ( Probability in favour ) / (Probability against)

= (3/8) / (5/8) = 3 / 5   Option B